OpenStudy (anonymous):

which of the following is geometric sequence 1) 1, 3, 5, 7,... 2) 1,3,6,10,... 3) 1, 0.1, 0.01, 0.001,... 4) 1, 1/2, 1/3, 1/4,...

jimthompson5910 (jim_thompson5910):

Let's check if 1, 3, 5, 7 is geometric

OpenStudy (anonymous):

how do u check it

jimthompson5910 (jim_thompson5910):

Divide the second term by the first term: 3/1 = 3 Then divide the third term by the second: 5/3 = 1.667 We get different results. If the sequence was geometric, then you'd get the same result everytime

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

Something like 1, 2, 4, 8, 16 is geometric because 2/1 = 2 4/2 = 2 8/4 = 2 16/8 = 2 Notice how I'm dividing each term by the previous term and I get the same result everytime: 2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

is 1,3,6,10,... geometric?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

why not

OpenStudy (anonymous):

because 10/6 is 1.6666...

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (anonymous):

6/3 is 1 and 3/1 is 1 so all of them has to be 1 in order to be geo sequence

jimthompson5910 (jim_thompson5910):

6/3 = 2 actually

jimthompson5910 (jim_thompson5910):

3/1 = 3

OpenStudy (anonymous):

oh right they have different number

jimthompson5910 (jim_thompson5910):

3/1 = 3 6/3 = 2 10/6 = 1.667 the results are NOT the same, so 1,3,6,10,... is NOT geometric

jimthompson5910 (jim_thompson5910):

again, I'm dividing each term by the previous term

OpenStudy (anonymous):

0.1/1=0.1 0.01/0.1=0.1 .001/0.01=0.1 so choice 3 is a geometric sequence because the result are the same

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

how about 1, 1/2, 1/3, 1/4

OpenStudy (anonymous):

(1/2)/1=.5 (1/3)/(1/2)=0.666666... not the same result

jimthompson5910 (jim_thompson5910):

nice job

jimthompson5910 (jim_thompson5910):

so only choice 3 is geometric

OpenStudy (anonymous):

thanks appreciate it

jimthompson5910 (jim_thompson5910):

you're welcome